FEL Upgrade Project Weekly Brief
February 19-23, 2001
 
 

Highlights:

The FEL operated well this week for the continuation of the winter FEL user run.  We were pleased to provide beam all
week to Prof. Robert Austin of Princeton university for continuation of his experiments on energy flow in biomolecules
using the amide I aborbance at 6 microns.

Management:

We completed the cost accounting for the January monthly report which will be distributed next week. As of Jan. 30,
2001 the project is 29% complete vs. 26% scheduled.  We have a small favorable schedule variance of +253k and a
small negative cost variance of -$228k.

We also completed a draft of a new cost performance plan that will accommodate the accumulated Ecru's during the
current Phase 1 of the project and the new Phase 2 of the project commencing this month. We plan to have this draft
reviewed by the contract monitor so that it can be used for the project cost performance reports starting with the
February report.

WBS 3 (Beam Physics):

A linac to linac UV bypass solution was completed. It provides for placement of the wiggler in the pit. Preliminary analysis
indicates single particle performance (chromatic, geometric) is adequate, but analysis is ongoing. Machine error, CSR, and other collective effects have not yet been investigated.

WBS 4 (Injector):

Continued work on field emission reduction. The FET system will be modified to collect more data, looked at
mechanisms involved in field emission reduction.

Gun HVPS - The vendor, Glassman, is ready to start work on the Gun HVPS as soon as we can get our spare unit shipped to them.

WBS 6 (RF):

Zone 3 - No Progress.  Actually, 4 Filament boards were removed and installed in zone 4.

Zone 4 - Four faulty Filament boards were replaced.  The problem of the Cathode Power Supply tripping during initial
turn on was solved by operating it in Zero Start By-Pass mode.  The vendor will be in next
week to tune the Circulator for klystron #8.  The discrepancy was resolved for the various RF power readings for
klystron #2.

Injector RF - CPI is having a problem getting their vendors deliver the various parts.  It appears the first 100 kW klystron
delivery will be about a month late at 5/18/01.  This will have no impact on the program.
There is nothing new to report on the upgrade of the HVPS for the Quarter.

WBS 8 (I&C):

This week continued with operations support. A divide-by-two was added to the 37 MHz monitor in the clean room to
produce a clean 18 MHz reference synchronous with the FEL micropulses, thanks to Hovator et.al.

New Printed circuit board for OBPMs is at the vendor.  Expect delivery early next week so we can get these
operational.  Thank you to EECAD for the rapid turn-around on the re-work. The new picomotor controller
chassis, which will control the transport mirrors, is being tested and should be ready soon.

Video problems in labs 2 and 3 have been corrected and we now have these areas available.  Work continues on the
user power supplies for each of the labs.

Four outstanding drawings were submitted to EECAD on 2/22 for the final updates and should be signed off soon.  We
are in the process of gathering existing drawings into a central location in the FEL for use
by everyone.

Discussions continue for commercial trim cards. A number of companies have expressed interest. The motivation for
changing is a higher current is needed and the replacement cost for the 10 amp cards
(complete 32 channel rack) is ~ $4000 per channel - my target price is <$2000. The new system would also address real
time monitoring  - the current system can not tell if the current had fluctuated putting you off loop.

A new FEL task management system is on the FEL website.  Anyone desiring to submit a task for optics or I & C are
encouraged to use this system. The link is: http://laser.jlab.org/task

WBS 9 (Transport):

Dipoles
Injector Dipoles (DU/DV)
o  DULY Research resolved their magnetic modelís anomalous tapered field results by going to a finer mesh.
    They continue working the on the 3D drafting model as well as the 2D details.
Optical Chicane Dipole (DW)
o  We continued incorporating backcheck information into the DW drawing package.
Arc Dipoles (GY, GX, GQ)
o  AES continued the layout the coil's leads and lead stabilization comb, incorporating the cover, the leads, the
    bus bars and the hose fittings into a single system.
o  AES continued the axisymmetric 2D magnetic model of the GY, finding there is indeed a 8.5 to 10 inch
    wide good field region over the full range of excitation using silicon steel boundaries that are within the coil
    on the inside edge and in the gap on the outside edge.  They are now creating a 3D magnetic model
    for final magnetic verification.
Quadrupoles
QX (3.125" Quad)
o  Measurement Probe: The coil support plateís coil slots are within acceptable tolerance so that the coils may
    be installed, the stabilization cap glued on and the end journals assembled.
o  Design Package: The QX drawing package was modified to incorporate the latest mounting and machining
    scheme and are in checking.
Sextupole with H/V Correctors
o  Robin Wines continued work on the magnetic model of the Sextupole, finding the automatic mesh
    generator makes meshes that yield unallowed harmonics, a clear indication of improper use of the model.
    She is working on creating a mesh that has boundaries that yield zero unallowed harmonics.
o  David Douglas clarified the specifications of the sextupole, significantly tightening them to 1 part per
    thousand over the good field region.  We will need the properly made mesh (above) to resolve errors of this
    level.
General
o  We reworked the budget such that the 180° bends, the optical chicane bends and the reverse bend
    chambers will be bought with Phase 1 funds.  This will have to increase the level of support from engineering
    for the chambers in order to achieve these critical procurements.
o The layout of the UV add-on was superposed on the overall machine layout and found to match so no
    changes to the IR machine were necessary.
o  Girder arrangements are firming up.  We will make individual girders for each quadrupole, re-use the old
    stands and cartridges, and use long beam pipes through the triplets, (forcing alignment to do a rough
    alignment before the tube is inserted).  The arrangements of diagnostics for the IR machine were
    simplified as the layout progressed.  The result was greater commonality of stands and girders leading to
    simplified design.
o We defined the alignment requirements for the dipoles with the help of Chris Curtis.  We will specify
    machined-in tooling ball holes at the factory with possible vendor CMM verification on all magnets
    other than the 180° Bends.  The latter, we will establish the fiducial locations on site.  Injector dipoles
    will have a surface for a machinistís level.

WBS 10 (Wiggler):

o Dispersion Section: The dispersion section is assembled and will be shipped next week.

WBS 11 (Optics):

We signed off on the drawing package for the deformable mirror assembly, and are awaiting results of our RFQ.    More
parts for the mirror test stand arrived, namely the flange that the electrical and water
feedthroughs are mounted on.  The vacuum vessel is being welded today, and will be leak checked early next week.

We ordered new quadrant pyroelectric detectors from Molectron, and will use them to see if we can get more reliable
performance over a larger range of displacements.  We've also determined that the electronics on the detector are very
pulsewidth dependent, and have asked I & C to implement a better scheme.  In the meantime, we have mounted the
PyroCAM to take the pick-off from the OBPM in the OCR (optics control room), and will use it to monitor the beam
position.

In the course of supporting operations, we supplied parts and technical assistance to our primary user, and assisted in
reconfiguring the hutch in User Lab 3 for metal-treatment research.  We had a brass fitting on one of our water-cooled
laser power meters split during owl shift, and when operations began in the morning, learned that some 7.5 gal of cooling
water had been dumped onto our purged optical table in the OCR.  We are working with the vendor to understand why
the fittings are splitting.

Operations/Commissioning:

See comments under highlights and in WBS 8 and 11 reports.



Privacy and Security Notice